OpenStudy (anonymous):

Having a little trouble? This is the problem I have to solve. Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y = 7/x^2 + 10 I used: f(x)=7/x g(x)=x^2+10 But when i plugged everything in I got 7/x^2+10. How can I fix that?

OpenStudy (anonymous):

close... let f(x) = 7/x + 10 and let g(x) = x^2

OpenStudy (amistre64):

is that +10 in the denominator?

OpenStudy (anonymous):

Yes, my answer had +10 in the denominator but it's not supposed to be.

OpenStudy (anonymous):

\[f(x) = \frac{ 7 }{ x }+10 \text{, }g(x)=x^2 \text{. Then }f(g(x)) = \frac{ 7 }{ g(x) }+10=\frac{ 7 }{ x^2 }+10\]

OpenStudy (amistre64):

another approach is: f(x) = x+10, g(x) = 7/x

OpenStudy (amistre64):

well, 7/x^2 lol

OpenStudy (anonymous):

yes but g(x) would have to be 7/x^2

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

I see where I was wrong now. Thank you guys so much.

OpenStudy (anonymous):

you're welcome!